OpenStudy (anonymous):

Simplify.. sqrt(8w^2 u^4) . sqrt(2w^9 u^6)

OpenStudy (anonymous):

any idea?

OpenStudy (anonymous):

Ammm maybe take term2 to the other side then we take the sqrt?

OpenStudy (anonymous):

then Squared the sqrt right

OpenStudy (anonymous):

first do thid \[\sqrt{term1} \times \sqrt{term2}=\sqrt{term1*term2}\]

OpenStudy (anonymous):

sqrt(16 w^11 u^11)

OpenStudy (anonymous):

u^10

OpenStudy (anonymous):

ok,can u do this now?

OpenStudy (anonymous):

1 min

OpenStudy (anonymous):

\[\sqrt{2^{4} \times w ^{11}\times u ^{10}}\]

OpenStudy (anonymous):

4w^11/2 u^5?

OpenStudy (anonymous):

okey now i have to take the sqrt right?

OpenStudy (anonymous):

ok,u r right =\[4u ^{5}w ^{5}\sqrt{w}\]

OpenStudy (anonymous):

Ahhh and can i write it like i did 4w^11/2 u^5?

OpenStudy (anonymous):

it is better to write \[w ^{5} \sqrt{w}\]

OpenStudy (anonymous):

okey thank u man (((:

OpenStudy (anonymous):

ur welcom :)